OpenStudy (anonymous):

One number is 6 more than twice another. If their sum is 51, find the numbers. Which of the following systems of equations represents the word problem?

OpenStudy (aroub):

What are the following?

OpenStudy (anonymous):

A. y = 2x + 6 and y = x + 51 B. = 2x + 6 and x + y = 51 C. y = 2(x + 6) and x + y = 51

OpenStudy (aroub):

B. y = 2x + 6 and x + y = 51